### v1

# Convergence of Random Products of Countably Infinitely Many Projections

Abstract.
Let $r\in\mathbb{N}\cup\{\infty\}$ be a fixed number and let $P_{j}\,\,(1\leq j\leq r)$ be the projection onto the closed subspace $\mathcal{M}_{j}$ of $\mathscr{H}$. We are interested in studying the sequence $P_{i_{1}},P_{i_{2}},\ldots\in\{P_{1},\ldots,P_{r}\}$. A significant problem is to demonstrate conditions under which the sequence $\{P_{i_{n}}\cdots P_{i_{2}}P_{i_{1}}x\}_{n=1}^{\infty}$ converges strongly or weakly to $Px$ for any $x\in\mathscr{H}$, where $P$ is the projection onto the intersection $\mathcal{M}=\mathcal{M}_{1}\cap\ldots\cap\mathcal{M}_{r}$. Several mathematicians have presented their insights on this matter since von Neumann established his result in the case of $r=2$. In this paper, we give an affirmative answer to a question posed by M. Sakai. We present a result concerning random products of countably infinitely many projections (the case $r=\infty$) incorporating the notion of pseudo-periodic function.
Key words and phrases:
Iterated sequence; random product of projections; pseudo-periodic function; strong convergence.
Mathematics Subject Classification:
46C05; 47A05; 47B02
1. Introduction
Throughout this note, let $\mathbb{B}(\mathscr{H})$ stand for the algebra of all bounded linear operators acting on a Hilbert space $(\mathscr{H},\langle\cdot,\cdot\rangle)$. The identity operator is denoted by $I$. By an (orthogonal) projection we mean an operator $P\in\mathbb{B}(\mathscr{H})$ such that $P^{2}=P=P^{*}$. As usual, $\mathscr{M}^{\perp}$ stands for the orthogonal complement of a (closed) subspace $\mathscr{M}$. The range and kernel of any operator $T$ are denoted by $\mathcal{R}(T)$ and $\mathcal{N}(T)$, respectively. In this context, we examine the convergence of a sequence $\{x_{n}\}$ in $\mathscr{H}$ strongly (that is, in norm) and weakly (that is, there exists some $x\in\mathscr{H}$ such that $\{\langle x_{n},y\rangle\}$ converges to $\langle x,y\rangle$ for all $y\in\mathscr{H}$).
Let $r\in\mathbb{N}\cup\{\infty\}$ be a fixed number and let $P_{j}\,\,(1\leq j\leq r)$ be the projection onto a closed subspace $\mathscr{M}_{j}$ of $\mathscr{H}$. Consider the sequence $P_{i_{1}},P_{i_{2}},\ldots\in\{P_{1},\ldots,P_{r}\}$. An interesting problem is posed as follows:
Problem. Under what conditions does the sequence $\{P_{i_{n}}\cdots P_{i_{2}}P_{i_{1}}x\}_{n=1}^{\infty}$ strongly or weakly converge to $Px$ for any vector $x\in\mathscr{H}$, where $P$ denotes the projection onto the intersection $\mathscr{M}=\mathscr{M}_{1}\cap\ldots\cap\mathscr{M}_{r}$?
For $r=3$, the sequence of iterates defined as $x_{n}=P_{i_{n}}x_{n-1}$ is illustrated in Figure 1.
Figure 1.
This problem is inherently complicated, in particular, in the case where $r=\infty$; refer to [6, 15].
Amemya and Ando [1] proved that the sequence $\{P_{i_{n}}\cdots P_{i_{2}}P_{i_{1}}x\}_{n=1}^{\infty}$ converges weakly to $Px$ for any vector $x\in\mathscr{H}$. They conjectured that their result is also valid if we replace “weakly” with “strongly".
For $r\geq 3$, the conjecture was already true if $\mathscr{H}$ is finite-dimensional; as shown in Práger [12].
Halperin [7, Theorem 1] established that if $Q_{i}$’s are chosen in such a way that $P_{i_{rn}}\cdots P_{i_{2}}P_{i_{1}}=(P_{r}\cdots P_{2}P_{1})^{n}$ for all $n$, then the subsequence $\{P_{i_{rn}}\cdots P_{i_{2}}P_{i_{1}}x\}_{n=1}^{\infty}$ of $\{P_{i_{n}}\cdots P_{i_{2}}P_{i_{1}}x\}_{n=1}^{\infty}$ strongly converges to $Px$. However, the convergence
$\lim_{n}\|(P_{1}\cdots P_{n})^{n}x-Px\|=0$
may not be uniform on bounded subsets of initial points $x$. Many mathematicians are working on the rate of convergence; see [2, 14] and references therein. Sakai [15] extended the Halperin result for quasi-periodic functions and a finite number of projections. He then posed a question of whether his results are still valid for a larger class than quasi-periodic functions or countably infinitely many projections. In [10], the authors addressed Sakai’s question for a broader class beyond quasi-periodic functions, which are called quasi-normal sequences; see [16]. In this paper, we explore the convergence of non-periodic products of projections, a topic which is also examined in various works like [13, 14].
The case $r=2$ was proved much earlier by von Neumann [17]. Indeed, he stated his well-known alternating projection theorem by showing that if $P_{1}$ and $P_{2}$ are projections, then the sequence $P_{1}x,P_{2}P_{1}x,P_{1}P_{2}P_{1}x,\ldots$ converges strongly to $Px$ in which $P$ is the projection onto the intersection of their corresponding closed subspaces.
However, Paszkiewicz [11] (with $r=5$) as well as Kopecká and V. Müller [8] (with $r=3$) gave counterexamples for the Amemiya–Ando conjecture. In general, the behavior of projections can be strange as shown by Kopecká and Paszkiewicz [9]. If $\mathscr{H}$ is an infinite-dimensional Hilbert space, then there exist three projections $P_{1},P_{2}$, and $P_{3}$ onto closed subspaces of $\mathscr{H}$ such that for every nonzero vector $x\in\mathscr{H}$, there exist integers $k_{1},k_{2},\ldots\in\{1,2,3\}$ such that $\{P_{k_{n}},\cdots P_{k_{2}}P_{k_{1}}x\}$ does not converge strongly.
Variations of this problem have applications in multiple disciplines such as partial differential equations, approximation theory, biomathematics, and computer science; see [4] and references therein.
In this paper, we provide a positive answer to Sakai’s question by presenting a result regarding random products of an infinite number of projections that involve a general notion of quasi-periodic functions, namely pseudo-periodic functions, which differ from quasi-normal sequences. We hope that this insight can contribute to solving the problem. We also provide concrete examples to clarify our results.
2. Main results
We start our work with the following notion appearing in [15].
Definition 2.1.
A map $\sigma:\mathbb{N}\to\{1,2,\ldots,r\}$ is called quasi-periodic if for each $1\leq j\leq r$, the inverse image of $\{j\}$ under $\sigma$ is an infinite set, and there is an integer $m\geq r$ such that for each integer $k\geq 0$, it holds that
$\{\sigma(k+1),\sigma(k+2),\ldots,\sigma(k+m)\}=\{1,2,\ldots,r\},$
or equivalently, if $\{l_{n}\}_{n=1}^{\infty}$ is the increasing sequence (depending on $j$) of all natural numbers such that $\sigma(l_{n})=j$ and $l_{0}=0$, then
$I(\sigma,j):=\sup_{n}\left(l_{n}-l_{n-1}\right)<\infty\,,\qquad(j\in\{1,\ldots,r\}).$
The following is an extension of Halperin’s result.
Theorem 2.2.
[15, Theorem] Let $r$ be a positive integer, let $P_{1},\ldots,P_{r}\in\mathbb{B}(\mathscr{H})$ be projections onto closed subspaces $\mathscr{M}_{1},\ldots,\mathscr{M}_{r}$ of $\mathscr{H}$, respectively, and let $P$ be the projection onto $\cap_{i=1}^{r}\mathscr{M}_{i}$ Let $\sigma:\mathbb{N}\to\{1,\ldots,r\}$ be quasi-periodic and set $T_{1}:=P_{\sigma(1)}$ and $T_{n}:=P_{\sigma(n)}T_{n-1}$. Then $\{T_{n}x\}$ strongy converges to $Px$. Furthermore,
$\|T_{n}x-T_{m}x\|^{2}\leq\big{(}(b-1)(b-2)+3\big{)}\sum_{k=m}^{n-1}\|T_{k+1}x-T_{k}x\|^{2}\,,$
where $x\in\mathscr{H}$, $b=\max_{1\leq j\leq r}I(\sigma,j)$ and $n>m\geq 1$.
Now, we extend the notion of quasi-periodic.
Suppose that $\sigma:\mathbb{N}\to\mathbb{N}$ is a random function such that for each $j\in\mathbb{N}$, the inverse image of $\{j\}$ under $\sigma$ is an infinite set, and let $\{l_{n}\}_{n=1}^{\infty}$ be the increasing sequence (depending on $j$) of all natural numbers such that $\sigma(l_{n})=j$ and $l_{0}=0$. Let $\Gamma_{F}$ denote the set of all $j$ such that
$I(\sigma,j)=\sup_{n}\left(l_{n}-l_{n-1}\right)<\infty.$ (2.1)
Set $\Gamma_{\infty}:=\mathbb{N}\backslash\Gamma_{F}$. Hence, $\Gamma_{F}$ and $\Gamma_{\infty}$ are disjoint subsets of $\mathbb{N}$ and $\Gamma_{F}\cup\Gamma_{\infty}=\mathbb{N}$.
Let $\{k_{n}\}$ denote the complement of the union of all sequences $\{l_{n}\}$ in the form of an increasing sequence. Evidently, $\sigma(k_{n})\in\Gamma_{\infty}$.
Definition 2.3.
A function $\sigma:\mathbb{N}\to\mathbb{N}$ is a pseudo-periodic function if $\Gamma_{F}=\{1,2,\ldots,r\}$ for some positive integer $r$ and $\{k_{n}-k_{n-1}\}$ is an increasing sequence. In this case, there exists an integer $m\geq r$ such that for each integer $k\geq 0$,
$\{1,2,\ldots,r\}\subseteq\{\sigma(k+1),\sigma(k+2),\ldots,\sigma(k+m)\}.$
The following is a typical example of random functions that we deal with.
Example 2.4.
Let $P_{1}$, $P_{2}$, and $P_{3}$ be arbitrary projections on a Hilbert space $\mathscr{H}$. Let $\{P_{i}\}_{i=4}^{\infty}$ be monotonically decreasing projections. Let $\sigma^{\prime}:\mathbb{N}\to\{4,5,\ldots\}$ be a random function. Define $\sigma:\mathbb{N}\to\mathbb{N}$ as follows:
$\sigma(n)=\begin{cases}1&n=3k-1\\ 2&n=3k-2\\ 3&n=3k,n\neq 3^{k^{\prime}}\\ \sigma^{\prime}(k^{\prime})&n=3k,n=3^{k^{\prime}}\\ \end{cases}\qquad\mbox{for~{}some~{} $k,k^{\prime}\in\mathbb{N}$}$
For example, when $j=3$, the sequence of indices for which $\sigma(l_{n})=3$ is $\{l_{n}\}=6,12,15,18,\ldots$.
It is easy to see that
$I(\sigma,1)=I(\sigma,2)=3\,,\qquad I(\sigma,3)=6\quad I(\sigma,4)=I(\sigma,5)=\ldots=\infty\,.$
This shows that $\sigma$ is not quasi-periodic. Here, we deal with the sequence
$T_{n}$ $:=P_{\sigma(n)}\cdots P_{\sigma(4)}P_{\sigma(3)}P_{\sigma(2)}P_{\sigma(1)}$
[…]
$=P_{\sigma(n)}\cdots P_{\sigma(k_{3})}P_{1}\cdots P_{3}P_{1}P_{2}P_{3}P_{1}P_{2}P_{3}P_{1}P_{2}P_{\sigma(k_{2})}P_{1}P_{2}P_{3}P_{1}P_{2}P_{\sigma(k_{1})}P_{1}P_{2}$
The sequence of $k_{1},k_{2},k_{3}\ldots$ is $3,9,27,\ldots$. Evidently, $\sigma$ is pseudo-periodic.
To achieve our result, we need several key lemmas.
Lemma 2.5.
Let $r$ be a positive integer and let $P_{1},\ldots,P_{r}\in\mathbb{B}(\mathscr{H})$ be projections onto closed subspaces $\mathscr{M}_{1},\ldots,\mathscr{M}_{r}$ of $\mathscr{H}$, respectively. Let $\sigma:\mathbb{N}\to\{1,\ldots,r\}$ be a random function and set $T_{1}:=P_{\sigma(1)}$ and $T_{n}:=P_{\sigma(n)}T_{n-1}$. Then for a fixed $k$, it holds that
$\lim_{n\to\infty}\|T_{n-k}x-T_{n}x\|=0.$
Proof.
The sequence $\{\|T_{n}x\|\}$ is decreasing since
$\|T_{n-1}x\|\geq\|T_{n}x\|\geq 0\qquad(n\geq 2).$
Hence,
$\{\|T_{n}x\|\}\mbox{~{}is a convergent sequence.}$. (2.2)
For each $n$, we have
$\|T_{n-1}x-T_{n}x\|^{2}$ $=\|(I-P_{\sigma(n)})T_{n-1}x\|^{2}$
$=\langle T_{n-1}x-P_{\sigma(n)}T_{n-1}x,T_{n-1}x-P_{\sigma(n)}T_{n-1}x\rangle$
$=\|T_{n-1}x\|^{2}-\|P_{\sigma(n)}T_{n-1}x\|^{2}$
$=\|T_{n-1}x\|^{2}-\|T_{n}x\|^{2}.$ (2.3)
Hence,
$\lim_{n\to\infty}\|T_{n-1}x-T_{n}x\|=0.$ (2.4)
By induction we get the result. ∎
Lemma 2.6.
With the notation of Definition 2.3, let $i\in\mathbb{N}$. Then for the unique integers $k_{n}$ and $k_{n+1}$ depending on $i$ such that $k_{n}\leq i<k_{n+1}$, it holds that
$\lim_{i\to\infty}\|T_{i}x-T_{k_{n+1}}x\|=\lim_{i\to\infty}\|T_{i}x-T_{k_{n}}x\|=0\,.$
Proof.
Consider the set $\mathscr{S}=\{T_{k_{n}+1}x,T_{k_{n}+2}x,\ldots,T_{k_{n+1}-1}x\}$. If we set $T_{k_{n}}x=y$, then
$\mathscr{S}=\{P_{k_{n}+1}y,P_{k_{n}+2}P_{k_{n}+1}y\,,\ldots\,,P_{k_{n+1}-1}\ldots P_{k_{n}+2}P_{k_{n}+1}y\}$
In fact, $\mathscr{S}$ is constructed from a part of a quasi-periodic sequence, since $\{k_{n}+1,k_{n}+2,\ldots,k_{n+1}-1\}\in\Gamma_{F}$. Hence, Theorem 2.2 ensures that
$\|T_{j}x-T_{i}x\|^{2}\leq M\sum_{k=i}^{j-1}\|T_{k+1}x-T_{k}x\|^{2},$ (2.5)
for a fixed number $M\geq 0$ and all $j$ satisfying $k_{n}\leq i<j\leq k_{n+1}$. From equality (2), we have
$\|T_{k+1}x-T_{k}x\|^{2}=\|T_{k}x\|^{2}-\|T_{k+1}x\|^{2}\,,$
for each $k$. Therefore, inequality (2.5) yields that
$\|T_{j}x-T_{i}x\|^{2}\leq M(\|T_{i}x\|^{2}-\|T_{j}x\|^{2})\,,\qquad(k_{n}\leq i<j\leq k_{n+1}).$
Take $j=k_{n+1}$. Utilizing the parallelogram law and Lemma 2.5, we have
$\lim_{i}\|T_{k_{n+1}}x-T_{i}x\|^{2}$ $\leq\lim_{i}2\|T_{k_{n+1}}x-T_{k_{n+1}-1}x\|^{2}+2\|T_{k_{n+1}-1}x-T_{i}x\|^{2}$
$\leq\lim_{i}2\|T_{k_{n+1}}x-T_{k_{n+1}-1}x\|^{2}+\lim_{i}2M(\|T_{i}x\|^{2}-\|T_{k_{n+1}-1}x\|^{2})$
$=0,$
Since as $i\to\infty$, we have $k_{n+1}\to\infty$. By the same reasoning we get
$\lim_{i\to\infty}\|T_{i}x-T_{k_{n}}x\|=0\,.$
The proof of the next lemma is straightforward.
Lemma 2.7.
Let $r$ be a positive integer and let $P_{1},\ldots,P_{r}\in\mathbb{B}(\mathscr{H})$ be projections onto closed subspaces $\mathscr{M}_{1},\ldots,\mathscr{M}_{r}$ of $\mathscr{H}$, respectively. Let $\{P_{i}\}_{i=r+1}^{\infty}$ be monotonically decreasing projections on $\{\mathscr{M}_{i}\}_{i=r+1}^{\infty}$. If $x=(I-P_{n+r}S)z$ for some $S\in\mathbb{B}(\mathscr{H})$ and $z\in\mathscr{H}$, then for each $m\geq n+r$ it holds that
$x=(I-P_{m}S)z\,.$
Proof.
Since $x=(I-P_{n+r}S)z$, we have
$P_{m}x=P_{m}z-P_{m}Sz\quad\mbox{and}\quad(I-P_{m})x=(I-P_{m})z.$
because $P_{m}P_{n+r}=P_{m}$. Hence,
$x$ $=P_{m}x+(I-P_{m})x=P_{m}z-P_{m}Sz+(I-P_{m})z=(I-P_{m}S)z.$
The next lemma is well-known in the literature, but we provide a proof for convenience.
Lemma 2.8.
Let $P_{1},\ldots,P_{n}\in\mathbb{B}(\mathscr{H})$ be (not necessarily distinct) projections onto closed subspaces $\mathscr{M}_{1},\ldots,\mathscr{M}_{n}$ of $\mathscr{H}$, respectively. Let $\tau$ be a permutation of the set $\{1,2,\ldots,n\}$ and let $T=P_{\tau(1)}P_{\tau(2)}\cdots P_{\tau(n)}$. Then
$\overline{\mathcal{R}(I-T)}\oplus\left(\cap_{i=1}^{n}\mathscr{M}_{i}\right)=\mathscr{H}.$
Proof.
It is enough to show that $\cap_{i=1}^{n}\mathscr{M}_{i}=\mathcal{N}(I-T^{*})$. If $x\in\cap_{i=1}^{n}\mathscr{M}_{i}$, then $P_{\tau(j)}x=x$ for all $1\leq j\leq n$. Hence, $P_{\tau(n)}\cdots P_{\tau(2)}P_{\tau(1)}x=x$. Thus, $x\in\mathcal{N}(I-T^{*})$.
To establish the converse, note that for any projection $Q\in\mathbb{B}(\mathscr{H})$, the equality $\|x\|^{2}=\|Qx\|^{2}+\|(I-Q)x\|^{2}$ ensures that if $\|Qx\|=\|x\|$, then $Qx=x$.
Let $x\in\mathcal{N}(I-T^{*})$. Then
$P_{\tau(n)}\cdots P_{\tau(2)}P_{\tau(1)}x=x$ (2.6)
and
$\|x\|=\|P_{\tau(n)}\cdots P_{\tau(2)}P_{\tau(1)}x\|\leq\|P_{\tau(n-1)}\cdots P_{\tau(2)}P_{\tau(1)}x\|\leq\|x\|,$
since the norm of any projection is not greater than $1$. Hence, $\|P_{\tau(n-1)}\cdots P_{\tau(2)}P_{\tau(1)}x\|=\|x\|$. By continuing this process, we arrive at $\|P_{\tau(1)}\|=\|x\|$. Therefore, $P_{\tau(1)}x=x$. Hence, (2.6) yields that $P_{\tau(n)}\cdots P_{\tau(2)}x=x$. Employing the same process, we conclude that $P_{\tau(j)}x=x$ for all $2\leq j\leq n$. Thus, $x\in\cap_{i=1}^{n}\mathscr{M}_{i}$. ∎
We establish our first result concerning the weak convergence of a random product of projections that involved a pseudo-periodic function.
Theorem 2.9.
Let $r$ be a positive integer and let $P_{1},\ldots,P_{r}\in\mathbb{B}(\mathscr{H})$ be distinct projections. Let $\{P_{i}\}_{i=r+1}^{\infty}$ be a monotonically decreasing sequence of (not necessarily distinct) projections. Let $\mathcal{R}(P_{i})=\mathscr{M}_{i}$ for each $i\in\mathbb{N}$.
Let $\sigma:\mathbb{N}\to\mathbb{N}$ be pseudo-periodic. Set $T_{1}:=P_{\sigma(1)}$ and $T_{n}:=P_{\sigma(n)}T_{n-1}$. Then the sequence $\{T_{n}x\}$ converges to $Px$ weakly, where $P$ is the projection onto $\cap_{i=1}^{\infty}\mathscr{M}_{i}$.
Proof.
Without loss of generality, we assume that $\Gamma_{F}=\{1,\ldots,r\}$, $\Gamma_{\infty}=\{r+1,r+2,\ldots\}$, and $P_{k_{n}}:=P_{n+r}$ for each $n\in\mathbb{N}$. With the notation in the definition of the pseudo-periodic function $\sigma$, $P_{\sigma(k_{n})}$ appears as the $k_{n}$-th projection from the right side in the definition of $T_{n}:=P_{\sigma(n)}\cdots P_{\sigma(4)}P_{\sigma(3)}P_{\sigma(2)}P_{\sigma(1)}$. Hence, we can write
$T_{k_{n}+m}$ $=P_{k_{n}+m}\ldots P_{k_{n}+1}P_{k_{n}}T_{k_{n}-1}$
$=P_{k_{n}+m}\ldots P_{k_{n}+1}P_{n+r}T_{k_{n}-1},$
where $m$ is the positive integer appears in definition of pseudo-periodic. Thus, $\{k_{n}+1,k_{n}+2,\ldots,k_{n}+m\}=\{1,2,\ldots,r\}$. Let $x,y\in\mathscr{H}$ be arbitrary. Utilizing Lemma 2.8 and assuming that $x=x_{1,n}+x_{2,n}$, where
$x_{1,n}\in\mathcal{R}(I-P_{n+r}P_{k_{n}+1}P_{k_{n}+2}\cdots P_{k_{n}+m})\quad\textrm{and}\quad x_{2,n}\in\cap_{i\in\Gamma_{F}}\mathscr{M}_{i}\cap\mathscr{M}_{n+r}$
we have
$\langle T_{k_{n}-1}y,x\rangle$ $=\langle T_{k_{n}-1}y,x_{1,n}+x_{2,n}\rangle$
$=\langle T_{k_{n}-1}y,(I-P_{n+r}P_{k_{n}+1}P_{k_{n}+2}\cdots P_{k_{n}+m})z_{n}\rangle+\langle T_{k_{n}-1}y,x_{2,n}\rangle,$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\mathrm{for~{}some~{}}z_{n}\in\mathscr{H}$
$=\langle(I-P_{n+r}P_{k_{n}+1}P_{k_{n}+2}\cdots P_{k_{n}+m})T_{k_{n}-1}y,z_{n}\rangle+\langle y,T_{k_{n}-1}x_{2,n}\rangle$
$=\langle(T_{k_{n}-1}-T_{k_{n}+m})y,z_{n}\rangle+\langle y,Px\rangle$
By Lemma 2.7, the sequence $\{z_{n}\}$ is a finite sequence, and hence bounded. Therefore, it is bounded. From Lemmma 2.5, we derive that $\{T_{k_{n}-1}x\}$ weakly converges to $Px$. Hence, $\{T_{k_{n}}x\}$ weakly converges to $Px$. For each $i$ there exists a positive integer $n$, depending on $i$, such that $k_{n}\leq i<k_{n+1}$.
It follows from the parallelogram law that
$\|T_{i}x-T_{k_{n}}x\|^{2}\leq 2\|T_{i}x-T_{k_{n+1}}x\|^{2}+2\|T_{k_{n+1}}x-T_{k_{n}}x\|^{2}$
Hence, by Lemma 2.6 and (2.4), we have $\lim_{i}\|T_{i}x-T_{k_{n}}x\|=0$. Therefore,
$\lim_{i}\langle T_{i}x,y\rangle=\langle T_{i}x-T_{k_{n}}x,y\rangle+\langle T_{k_{n}}x,y\rangle=\langle Px,y\rangle\,.$
This shows that $\{T_{i}x\}$ weakly converges to $Px$.
Additional conditions are required to guarantee the strong convergence of the sequence ${T_{n}x}$ to $Px$. Specific terminology and lemmas are needed to achieve this result.
The following notion is introduced in [3].
Definition 2.10.
The angle of an $r$-tuple of closed subspaces $(\mathscr{M}_{1},\mathscr{M}_{2},\ldots,\mathscr{M}_{r})$ in a Hilbert space $\mathscr{H}$ is the angle in $[0,\frac{\pi}{2}]$ whose cosine is defined as
$c_{b}(\mathscr{M}_{1},\mathscr{M}_{2},\ldots,\mathscr{M}_{r})=\|P_{r}\ldots P_{1}P_{(\cap_{i=1}^{r}\mathscr{M}_{i})^{\perp}}\|$
The following lemma is interesting on its own right.
Lemma 2.11.
[3]
$c_{b}(\mathscr{M}_{1},\mathscr{M}_{2},\ldots,\mathscr{M}_{r})<1$
if and only if
$\mathscr{M}_{1}^{\perp}+\mathscr{M}_{2}^{\perp}\ldots+\mathscr{M}_{r}^{\perp}\,.$
is closed.
Remark 2.12.
In view of the above lemma, if $c_{b}(\mathscr{M}_{1},\mathscr{M}_{2},\ldots,\mathscr{M}_{r})<1$, then for each $t>r$ we have
$c_{b}(\mathscr{N}_{1},\mathscr{N}_{2},\ldots,\mathscr{N}_{t})<1\,,$
when $\{\mathscr{N}_{i}:1\leq i\leq t\}=\{\mathscr{M}_{i}:1\leq i\leq r\}$.
The proof of the subsequent lemma is straightforward; therefore, we omit it.
Lemma 2.13.
Let $Q$ be the projection onto a closed subspace of a Hilbert space. Let $x$ and $y$ be elements in the space. Then
$\|x-y\|^{2}\leq\|x-Qy\|^{2}+\|x-Qx\|^{2}+2\|y-Qy\|^{2}.$ (2.7)
The next lemma read as follows.
Lemma 2.14.
Let $\{P_{i}\}_{i\in J}$ be finite or infinite Projection on Hilbert space $\mathscr{H}$. Let $\sigma:\mathbb{N}\to J$ be a random function. Set $T_{1}:=P_{\sigma(1)}$ and $T_{n}:=P_{\sigma(n)}T_{n-1}$. Let $T_{n}x$ is weakly converges to $Px$, where $P$ is the projection onto $\cap_{i\in J}\mathcal{R}(P_{i})$. If there is a subsequence $T_{n_{k}}$ such that converges to $Px$ strongly, then $T_{n}$ converges to $Px$ strongly.
Proof.
Since
$\|T_{n-1}x\|^{2}\geq\|T_{n}x\|^{2}\,,\quad(x\in\mathscr{H})\,,$
$|T_{n-1}|^{2}\geq|T_{n}|^{2}$, and hence $|T_{n-1}|\geq|T_{n}|$. Then by Vigier Theorem $|T_{n}|$ converges to an operator $S$ strongly. Thus,
$\lim_{n\to\infty}\|T_{n}x\|=\|Sx\|\,.$ (2.8)
On the other hand, $\lim_{n\to\infty}|T_{n_{k}}x|=\|Px\|$, which gives that $\|Px\|=\|Sx\|$. Now weakly convergence of $T_{n}x$ and (2.8) completes the proof. ∎
Our next main result is as follows.
Theorem 2.15.
Let $r$ be a positive integer and let $P_{1},\ldots,P_{r}\in\mathbb{B}(\mathscr{H})$ be distinct projections such that $c_{b}(\mathcal{R}(P_{1}),\mathcal{R}(P_{2}),\mathcal{R}(P_{r}))<1$. Let $\{P_{i}\}_{i=r+1}^{\infty}$ be a monotonically decreasing sequence of (not necessarily distinct) projections such that $\cap_{i=1}^{r}\mathcal{R}(P_{i})=\cap_{i=1}^{\infty}\mathcal{R}(P_{i})$. Let $\sigma:\mathbb{N}\to\mathbb{N}$ be pseudo-periodic. Set $T_{1}:=P_{\sigma(1)}$ and $T_{n}:=P_{\sigma(n)}T_{n-1}$. Then the sequence $\{T_{n}x\}$ converges to $Px$ strongly, where $P$ is the projection onto $\cap_{i=1}^{\infty}\mathcal{R}(P_{i})$.
Proof.
Let $x\in\mathscr{H}$. It follows from Theorem 2.9 that $T_{n}x$ weakly converges to $Px$. Fix $i_{0}<j_{0}$ be positive integers. For each positive integer $i$ with $i_{0}\leq i\leq j_{0}$ and each positive integer $t$ (depending on $i$) with $k_{i}\leq t\leq k_{i+1}-1$, Lemma 2.13 with $Q=P_{t+1}$ ensures that
$\|Px-T_{t}x\|^{2}\leq\|Px-T_{t+1}x\|^{2}+2\|T_{t}x-T_{t+1}x\|^{2},$ (2.9)
since $P_{j}P=PP_{j}=P$ for all $j$. By using Lemma 2.13 with $Q=P_{k_{i+1}}$, we can further deduce that
$\|Px-T_{k_{i+1}-1}x\|^{2}$ $\leq\|Px-T_{k_{i+1}}x\|^{2}+\|Px-P_{k_{i+1}}Px\|^{2}$
$+2\|T_{k_{i+1}-1}x-T_{k_{i+1}}x\|^{2}$
$=\|Px-T_{k_{i+1}}x\|^{2}+2\|T_{k_{i+1}-1}x-T_{k_{i+1}}x\|^{2}\,.$ (2.10)
By adding inequalities (2.9) over all $k_{i}\leq t\leq k_{i+1}-2$ and adding (2), we have
$\|Px-T_{k_{i}}x\|^{2}\leq\|Px-T_{k_{i+1}}x\|^{2}+2\sum_{t=k_{i}}^{k_{i+1}-1}\|T_{t}x-T_{t+1}x\|^{2}$ (2.11)
for each fixed $i_{0}\leq i\leq j_{0}$. By adding inequalities (2.11) over all $i_{0}\leq i\leq j_{0}$, we reach
$\|Px-T_{k_{i_{0}}}x\|^{2}$ $\leq\|Px-T_{k_{j_{0}+1}}x\|^{2}+2\sum_{t=k_{i_{0}}}^{k_{{j_{0}+1}}-1}\|T_{t}x-T_{t+1}x\|^{2}$
$\leq\|Px-T_{k_{j_{0}+1}}x\|^{2}+2\sum_{t=k_{i_{0}}}^{k_{{j_{0}+1}}-1}(\|T_{t}x\|^{2}-\|T_{t+1}x\|^{2})$
$=\|Px-T_{k_{j_{0}+1}}x\|^{2}+2(\|T_{k_{i_{0}}}x\|^{2}-\|T_{k_{j_{0}+1}}x\|^{2})$
$\leq 2\|Px-T_{k_{j_{0}+1}-1}x\|^{2}+2\|T_{k_{j_{0}+1}}x-T_{k_{j_{0}+1}-1}x\|^{2}$
$\quad+2(\|T_{k_{i_{0}}}x\|^{2}-\|T_{k_{j_{0}}}x\|^{2}),$ (2.12)
where we use the parallelogram law to get the last inequality.
Note that $\|P_{r}\cdots P_{1}(I-P)\|<1$ and $Q_{1}\cdots Q_{m}$ is a product of $m\geq r$ projections such that $\{Q_{1},\ldots,Q_{m}\}=\{P_{1},\ldots,P_{r}\}$. Remark 2.12 implies that $\|Q_{1}\cdots Q_{m}(I-P)\|<1$. Hence, there exists a constant $c<1$ such that $\|P_{\sigma(n+m)}\ldots P_{\sigma(n+1)}(I-P)\|<c$ whenever $k_{j_{0}}<n+1<\cdots<n+m<k_{j_{0}+1}$. Therefore,
$\|Px-T_{k_{j_{0}+1}-1}x\|$ $=\|PT_{k_{j_{0}}}x-T_{k_{j_{0}+1}-1}x\|$
$\leq\|(P_{k_{j_{0}+1}-1}P_{k_{j_{0}+1}-2}\cdots P_{k_{j_{0}}+1}-P)T_{k_{j_{0}}}x\|$
$=\|(P_{k_{j_{0}+1}-1}P_{k_{j_{0}+1}-2}\cdots P_{k_{j_{0}}+1}(1-P))T_{k_{j_{0}}}x\|$
$\leq\|P_{k_{j_{0}+1}-1}P_{k_{j_{0}+1}-2}\cdots P_{k_{j_{0}+1}-m}(1-P)\|$
×$\quad\times\|P_{k_{j_{0}+1}-m-1}\cdots P_{k_{j_{0}}-2m}(1-P)\|\ldots$
×$\quad\times\|P_{k_{j_{0}}+s+2m}\cdots P_{k_{j_{0}}+s+m+1}(1-P)\|\qquad(0\leq s\leq m-1)$
×$\quad\times\|P_{k_{j_{0}}+m+s}\cdots P_{k_{j_{0}}+1}\|\|T_{k_{j_{0}}}x\|$
$\leq c^{\nu}\|x\|,$
for some $\nu$, where $m$ is given in the Definition of quasi-periodic. Since $\{k_{n}-k_{n-1}\}$ is an increasing sequence, $\nu\to\infty$, if $j_{0}\to\infty$. Hence, (2.2), (2.4), and (2) ∎
The next example illustrates Theorem 2.15.
Example 2.16.
Let $\mathscr{H}$ be a separable Hilbert space with $\{e_{i}:i\in\mathbb{N}\}$ as its orthonormal basis. Let
$\mathscr{M}_{1}=\mathrm{span}\{e_{2k-1}:k\in\mathbb{N}\},\mathscr{M}_{2}=\mathrm{span}\left\{\frac{e_{2k-1}+e_{2k}}{2}:k\in\mathbb{N}\right\},\mathscr{M}_{i}=\mathrm{span}\{e_{3j}:j\geq i-2\}\,,$
for $i\geq 3$. Let $P_{i}$ be the projection onto $\mathscr{M}_{i}$ for each $i\geq 1$. It is easy to verify that $\mathscr{M}_{1}+\mathscr{M}_{2}=\mathscr{H}$ is closed. Therefore, $\mathscr{M}_{1}^{\perp}+\mathscr{M}_{2}^{\perp}=(\mathscr{M}_{1}\cap\mathscr{M}_{2})^{\perp}=\mathscr{H}$ is closed (see [5, Lemma 11]). In addition, $\mathscr{M}_{1}\cap\mathscr{M}_{2}=\{0\}=\cap_{i=3}^{\infty}\mathscr{M}_{i}$. It follows from Lemma 2.11 that
$c_{b}(\mathscr{M}_{1},\mathscr{M}_{2})<1\,.$
Let $P_{i}$ be the projection onto $\mathscr{M}_{i}$ for each $i\geq 1$. Let $\sigma:\mathbb{N}\to\mathbb{N}$ be pseudo-periodic with $\Gamma_{F}=\{1,2\}$. Set $T_{1}:=P_{\sigma(1)}$ and $T_{n}:=P_{\sigma(n)}T_{n-1}$. From Theorem 2.15 we conclude that the sequence $\{T_{n}x\}$ converges to 00 strongly.
Disclosure statement
On behalf of the authors, the corresponding author states that there is no conflict of interest. Data sharing is not applicable to this paper as no datasets were generated or analysed during the current study.